OpenStudy (anonymous):

Given that sin(A+B)=4/5 and cos(A-B)=12/13 with the condition that 0

OpenStudy (anonymous):

Okay ..  => tan(A+B) = 4/3 And tan(A-B) = 5/12 tan{(A+B)+(A-B)} = tan(A+A) = tan(2A) tan{(A+B)+(A-B)} \[\frac{\tan(A+B) + \tan(A-B)}{1- \tan(A+B) \tan(A-B)}\] Substitute the values and you'd get the answer :D